OpenStudy (shesolitt):

I WILL MEAL AND FAN Joe earns $40 for playing music on Saturday nights at the local coffee house. He makes an average of $3.25 in tips per hour. Write the function of Joe's earnings, and solve for how much he would make after 5 hours. y = 3.25x + 40; $43.25 y = 3.25x + 40; $56.25 y = 40x + 3.25; $58.25 y = 40x + 3.25; $203.25

OpenStudy (shesolitt):

my answer is B @Will.H

OpenStudy (will.h):

He makes 3.25 per hour so that means 3.25x and he makes an addition of 40 on Saturday nights so Y = 3.25x + 40 Solve when x = 5 Y = 3.25 * 5 + 40 Y = 56.25 So yes B

OpenStudy (shesolitt):

<3

OpenStudy (will.h):

:)